Solve.
sqrt( 7x+46 ) = x + 4

Answer by checkley71(8403)

sqrt(7x+46)=x+4
square both sides:
7x+46=x^2+8x+16
x^2+8x-7x+16-46=0
x^2+x-30=0
(x+6)(x-5)=0
x+6=0
x=-6 answer.
x-5=0
x=5 answer.